OpenStudy (anonymous):

Solve for x. log6(2x + 3) = 3

OpenStudy (campbell_st):

raise both sides to the power of 6 and remember 6^log6(2x + 3) = 2x+3 so 2x + 3 = 6^3 2x + 3 = 216 2x = 213 x = 106.5

OpenStudy (anonymous):

is it... \[\huge \log_{6}(2x+3)=3\] or \[\huge \log6(2x+3)=3\] ?

OpenStudy (anonymous):

\[\log _{6}(2x+3)=3\]